# Hybrid quantum–classical matrix-product state and Lanczos methods for electron–phonon systems with strong electronic correlations: Application to disordered systems coupled to Einstein phonons

### v1

@@ III.2 Lanczos-MTE method
Figure 2: Comparison of TEBD-MTE ($\delta=10^{-8}$}), Lanczos-MTE ($\epsilon=10^{-9}/t_{\text{f}}$), and regular MTE as implemented in Ref. [TenBrink2022] for a non-interacting system ($L=13$, $V=0$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\gamma=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$, $W=0$). Lanczos-MTE and TEBD use $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$ and all methods are averaged over $N_{\mathrm{traj}}=4000$ trajectories. The inset shows the difference $\delta\mathcal{O}_{\mathrm{CDW}}$ between the many-body hybrid techniques and the reference data from regular MTE from Ref. [TenBrink2022].
@@ III.2.1 Algorithm
The accuracy of the Lanczos step is controlled both by increasing the dimension of the constructed Krylov subspace and by reducing the Lanczos time-step $\Delta t_{\mathrm{Lanczos}}$ [Manmana2005]. As there exist readily available measurements for the error acquired in the time-dependent Lanczos step we can use an adaptive algorithm to fix the rate of error $\epsilon$ in the time-dependent Lanczos method, to an arbitrary value. Choosing the Krylov dimension and $\Delta t_{\mathrm{Lanczos}}$ adaptively means that we can combine the algorithmic parameters for the time-dependent Lanczos method into the single parameter $\epsilon$.
Note that whenever the Lanczos step has to be decreased to ensure accuracy, resulting in $\Delta t_{\mathrm{Lanczos}}<\Delta t$, we still want to fix $\Delta t$ separately which is given by the alternating evolution of environment and subsystem in Eq. 19. In our concrete simulations, however, it is sufficient to allow for a maximum Krylov dimension $\approx 30$ to ensure that the time-evolution is accurate without a need to reduce $\Delta t_{\mathrm{Lanczos}}$; thus $\Delta t_{\mathrm{Lanczos}}=\Delta t$ throughout this work. Our implementation of time-dependent Lanczos is based on [Haegeman2024].
@@ III.3 Benchmark
To demonstrate the performance of the Lanczos-MTE method we begin by comparing our results to those from Ref. [TenBrink2022] for a non-interacting system, which is shown in Fig. 2. While in this case the less expensive single-particle formalism can be employed to evolve the quantum system, we show data that has been generated using the many-body Lanczos-MTE method. For the example of $\mathcal{O}_{\mathrm{CDW}}$ we show that Lanczos-MTE and MTE produce equivalent results. The inset in Fig. 2 shows how the difference $\delta\mathcal{O}_{\mathrm{CDW}}$ between the different methods saturates around $\delta\mathcal{O}_{\mathrm{CDW}}\lesssim 10^{-2}$. This is indeed expected as we consider an ensemble of $N_{\mathrm{traj}}=4000$ trajectories: Based on the statistical convergence analysis from [TenBrink2022] carried out for the same system size, the accuracy of observables in the ensemble, such as $\mathcal{O}_{\mathrm{CDW}}$, is $\propto 1/\sqrt{N_{\mathrm{traj}}}$. The range of the imbalance $\mathcal{O}_{\mathrm{CDW}}(t)$ is [1,1][-1,1] and therefore, we can upper bound the statistical accuracy of $\mathcal{O}_{\mathrm{CDW}}(t)$ (and by extension also $\delta\mathcal{O}_{\mathrm{CDW}}(t)$) to be $2/\sqrt{N_{\mathrm{traj}}}=10^{-2}$.
@@
In order to asses the accuracy of the Lanczos-MTE method we first analyze its dependence on $\epsilon$. While the convergence behavior of the Lanczos step is generally well understood [Manmana2005], we want to show how the accuracy of the Lanczos step influences the ensemble convergence in the MTE method. In Fig. 3 we show the deviation $\delta\mathcal{O}_{\mathrm{CDW}}=|\mathcal{O}_{\mathrm{CDW},\epsilon^{*}}-\mathcal{O}_{\mathrm{CDW},\epsilon}|$ where $\epsilon^{*}=10^{-10}/t_{f}$. We observe that, again, the deviation $\delta\mathcal{O}_{\mathrm{CDW}}$ is never larger than $10^{-2}$. We conclude that for all considered values of $\epsilon$ the deviation between the compared ensemble is dominated by the statistical error, indicating that Lanczos-MTE is well behaved with respect to the error rate $\epsilon$.
@@
For local Hilbert spaces of dimension $2$, pure-state techniques such as Lanczos can reach system sizes as large as $L\simeq 30$ sites (see, e.g., [Steinigeweg2014, Herbrych2012]). In combination with MTE, the sampling over trajectories requires performing such calculations $N_{\mathrm{traj}}$ times, which adds computational costs that scale linearly in $N_{\mathrm{traj}}$.
@@ III.4 TEBD-MTE method
Similar to the Lanczos-MTE discussed in the previous section (Sec. III.2), we can use tensor network-based methods, such as the time-evolving block decimation (TEBD) method, for the quantum time evolution in Eq. 17. The evolution of the classical coordinates and momenta in Eq. 17 is done in the same way as before. This combined TEBD-MTE method brings all the advantages of the TEBD method, such as simulating larger systems compared to the Lanczos-MTE and may allow applications to equilibrium and dynamical quantities. In Fig. 2, we have compared both the Lanczos-MTE and the TEBD-MTE method, which agree with each other within the error limits mentioned before. Our TEBD implemtation utilizes the ITensor library [itensor].
@@ III.4.1 Algorithm
In TEBD-MTE, the quantum many-body state $|\psi(t)\rangle$ of the electronic sub-system is described by a matrix product state (MPS). The combined state of the whole system $(|\psi_{\{x_{\ell},p_{\ell}\}}(t)\rangle,\{x_{\ell},~p_{\ell}\})$ is graphically represented in the bottom layer of Fig. 1. Figure 1 describes one graph element associated with a site $\ell$, which contains information about both the sub-systems, classical and quantum. Here, $A_{a_{\ell-1},a_{\ell}}^{\sigma_{\ell}}$ are the matrices of the MPS where $a_{\ell-1},a_{\ell}$ are the bond indices, and $\sigma_{\ell}$ is the physical index [Schollwoeck2011] and $x_{\ell}$ and $p_{\ell}$ represent the position and momentum coordinates of the classical phonon on site $\ell$.
@@
Figure 4: Dependence of the TEBD-MTE method on the discarded weight $\delta$ and the time step $\Delta t$ in an interacting system ($L=14$, $V=$2\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\gamma=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$, $W=0$). We plot the deviation $\delta\mathcal{O}_{\mathrm{CDW}}$ when 4 varying $\delta$ at a fixed time step $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$, comparing with a reference set using $\delta=10^{-8}$ and 4 varying $\Delta t$ at fixed cutoff $\delta=10^{-8}$}, comparing with reference set using $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$}. The results are averaged over $N_{\mathrm{traj}}=4000$ trajectories.
@@
As we know [Schollwoeck2011, Paeckel2019], in the TEBD method, the error builds up with time, which depends on the time step in the Trotter-Suzuki decomposition and the truncation of the singular values during the singular value decomposition (SVD). The latter is related to the entanglement increase in a nonequilibrium problem and leads to an increase in the required resources. We can control the error of the TEBD step by decreasing the time step and the discarded weight.
@@ III.5 Benchmark
We envision several interesting applications of TEBD-MTE beyond the cases covered here. Generally, one of the most important applications of time-dependent MPS methods is the calculation of equilibrium spectral functions and other dynamical properties. These are usually cases with a mild entanglement increase, even at finite temperature. While the regime of intermediate to large values of $\omega_{0}$ have been covered with full-quantum MPS methods [Jansen2020, Jansen2023], the adiabatic regime of optical phonons with interacting electrons is mostly unexplored with MPS methods, which also applies to accustic phonons. In the long run, using hybrid MPS-trajectory-based methods in DMFT might be an interesting direction as well. In order to illustrate that TEBD-MTE is well behaved concerning its convergence properties on larger systems, we present further results in Appendix B.
@@ IV Stability of an Anderson insulator
In this section, we discuss the case of non-interacting electrons ($V=0$) with finite disorder. When the electronic sub-system is isolated from the phonons, the system exhibits Anderson localization for any finite disorder strength in one dimension [Kramer1993]. In this section we study the stability of the Anderson localization in the presence of classical optical phonons.
We want to highlight that we expect that for small $\gamma$ and $\omega_{0}$, MTE will accurately reproduce the electronic observables even of the original quantum model. However, we will go beyond the regime of $\unit\gamma\ll\unit{}$ addressing the question of the stability of Anderson laocalization against a classical bath for arbitrary values of $\gamma$, regardless of whether the quantum version is accurately described or not. The same philosophy will apply in Sec. V. This section serves to set the stage for the analysis of systems with electronic interactions; so long as $V=0$, no many-body treatment of the electronic sector is necessary and instead, we use single-particle orbitals in the MTE implementation as in [TenBrink2022]. We first study the spreading of a single electron and then move on to the decay of the CDW initial state.
@@ IV.1 Spreading of a single electron
Figure 7: Time-evolved density profile $\langle\hat{n}_{\ell}(t)\rangle$ of a particle initially ($t=0$) localized on a single lattice site $\ell=50$ on a lattice with $L=100$ sites and with disorder ($W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$}). We show the 7 uncoupled case $\gamma=0$, exhibiting Anderson localization, and 7 a system with a large electron–phonon coupling $\unit\gamma=\unit{}$. Further parameters are $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$ and the data was generated using the Lanczos-MTE routine with $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$, $\epsilon=10^{-9}/t_{f}$, and $N_{\mathrm{dis}}=5000$ realizations for $N_{\mathrm{traj}}=200$ trajectories each.
@@
Figure 8: Spreading of a single particle in the disordered system coupled to classical phonons. For a few values of $\gamma$ we show in 8 the MSD for the particle as a function of time and in the inset the time-dependent dynamical exponent $\bar{z}(t)$, averaged over a time window of $t^{*}=$800\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$ (see Eq. 30). In 8 the extracted estimates for the dynamical exponent $z$ is shown. The model parameters are $L=100$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, and $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$ and the particle is initially localized at lattice site $\ell=50$. The data are generated using an MTE routine using exact diagonalization in the single-particle picture to propagate the quantum-state with $N_{\mathrm{dis}}=2000$ realizations for $N_{\mathrm{traj}}=200$ trajectories each and $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$.
@@
Figure 9: Similar setup to Fig. 8 with $\unit\gamma=\unit{}$ held fixed while $W$ is varied. We show the dependency of the dynamical exponent $z$ on disorder strength obtained from MTE. Our estimate of the dynamical exponent $z$ is averaged over a large time window $t^{*}=$800\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$. The solid symbols are data from a full quantum-treatment of electrons coupled to dispersive hard-core bosons for $\unit\omega_{0}=\gamma=\unit{}$ from [Prelovsek2018], with a hard-core boson hopping amplitude $t_{b}=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$.
Anderson localization implies no transport at all in one dimension [Evers2008], so any persistent spreading of a particle can be interpreted as a breakdown of Anderson localization. In Fig. 7 we show how a particle on a lattice of $L=100$ sites and disorder strength $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$ spreads out in the system over time. The particle is initially localized at site $\ell=50$. For $\gamma=0$ we clearly see that the particle stays localized as expected. In contrast, when $\unit\gamma=\unit{}$, the particle continuously spreads out with a characteristic cone whose shape is related to the specific transport properties of the system.
@@
Interestingly, the qualitative and quantitative dependence of $z$ on $W$ is very similar to that obtained for $\unit\omega_{0}=\gamma=\unit{}$ obtained in [Prelovsek2018] from a numerically exact approach studying an electronic system coupled to dispersive hard-core bosons (see the circles in Fig. 9).
In summary, we observe delocalization and subdiffusive electronic transport due to the coupling to classical phonons. Going to more electrons and interacting system is expected to accelerate that delocalization which we will capture by comparing the decay of $\mathcal{O}_{\text{CDW}}$.

### v2

@@ III.2 Lanczos-MTE method
Figure 2: Comparison of TEBD-MTE ($\delta=10^{-8}$}), Lanczos-MTE ($\epsilon=10^{-9}/t_{\text{f}}$), and regular MTE as implemented in Ref. [97] for a non-interacting system ($L=13$, $V=0$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\gamma=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$, $W=0$). Lanczos-MTE and TEBD use $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$ and all methods are averaged over $N_{\mathrm{traj}}=4000$ trajectories. The inset shows the difference $\delta\mathcal{O}_{\mathrm{CDW}}$ between the many-body hybrid techniques and the reference data from regular MTE from Ref. [97].
@@ III.2.1 Algorithm
The accuracy of the Lanczos step is controlled both by increasing the dimension of the constructed Krylov subspace and by reducing the Lanczos time-step $\Delta t_{\mathrm{Lanczos}}$ [64]. As there exist readily available measurements for the error acquired in the time-dependent Lanczos step we can use an adaptive algorithm to fix the rate of error $\epsilon$ in the time-dependent Lanczos method, to an arbitrary value. Choosing the Krylov dimension and $\Delta t_{\mathrm{Lanczos}}$ adaptively means that we can combine the algorithmic parameters for the time-dependent Lanczos method into the single parameter $\epsilon$.
Note that whenever the Lanczos step has to be decreased to ensure accuracy, resulting in $\Delta t_{\mathrm{Lanczos}}<\Delta t$, we still want to fix $\Delta t$ separately which is given by the alternating evolution of environment and subsystem in Eq. 19. In our concrete simulations, however, it is sufficient to allow for a maximum Krylov dimension $\approx 30$ to ensure that the time-evolution is accurate without a need to reduce $\Delta t_{\mathrm{Lanczos}}$; thus $\Delta t_{\mathrm{Lanczos}}=\Delta t$ throughout this work. Our implementation of time-dependent Lanczos is based on [35].
@@ III.3 Benchmark
To demonstrate the performance of the Lanczos-MTE method we begin by comparing our results to those from Ref. [97] for a non-interacting system, which is shown in Fig. 2. While in this case the less expensive single-particle formalism can be employed to evolve the quantum system, we show data that has been generated using the many-body Lanczos-MTE method. For the example of $\mathcal{O}_{\mathrm{CDW}}$ we show that Lanczos-MTE and MTE produce equivalent results. The inset in Fig. 2 shows how the difference $\delta\mathcal{O}_{\mathrm{CDW}}$ between the different methods saturates around $\delta\mathcal{O}_{\mathrm{CDW}}\lesssim 10^{-2}$. This is indeed expected as we consider an ensemble of $N_{\mathrm{traj}}=4000$ trajectories: Based on the statistical convergence analysis from [97] carried out for the same system size, the accuracy of observables in the ensemble, such as $\mathcal{O}_{\mathrm{CDW}}$, is $\propto 1/\sqrt{N_{\mathrm{traj}}}$. The range of the imbalance $\mathcal{O}_{\mathrm{CDW}}(t)$ is [1,1][-1,1] and therefore, we can upper bound the statistical accuracy of $\mathcal{O}_{\mathrm{CDW}}(t)$ (and by extension also $\delta\mathcal{O}_{\mathrm{CDW}}(t)$) to be $2/\sqrt{N_{\mathrm{traj}}}=10^{-2}$.
@@
In order to asses the accuracy of the Lanczos-MTE method we first analyze its dependence on $\epsilon$. While the convergence behavior of the Lanczos step is generally well understood [64], we want to show how the accuracy of the Lanczos step influences the ensemble convergence in the MTE method. In Fig. 3 we show the deviation $\delta\mathcal{O}_{\mathrm{CDW}}=|\mathcal{O}_{\mathrm{CDW},\epsilon^{*}}-\mathcal{O}_{\mathrm{CDW},\epsilon}|$ where $\epsilon^{*}=10^{-10}/t_{f}$. We observe that, again, the deviation $\delta\mathcal{O}_{\mathrm{CDW}}$ is never larger than $10^{-2}$. We conclude that for all considered values of $\epsilon$ the deviation between the compared ensemble is dominated by the statistical error, indicating that Lanczos-MTE is well behaved with respect to the error rate $\epsilon$.
@@
For local Hilbert spaces of dimension 22, pure-state techniques such as Lanczos can reach system sizes as large as $L\simeq 30$ sites (see, e.g., [93, 37]). In combination with MTE, the sampling over trajectories requires performing such calculations $N_{\mathrm{traj}}$ times, which adds computational costs that scale linearly in $N_{\mathrm{traj}}$.
@@ III.4 TEBD-MTE method
Similar to the Lanczos-MTE discussed in the previous section (Sec. III.2), we can use tensor network-based methods, such as the time-evolving block decimation (TEBD) method, for the quantum time evolution in Eq. 17. The evolution of the classical coordinates and momenta in Eq. 17 is done in the same way as before. This combined TEBD-MTE method brings all the advantages of the TEBD method, such as simulating larger systems compared to the Lanczos-MTE and may allow applications to equilibrium and dynamical quantities. In Fig. 2, we have compared both the Lanczos-MTE and the TEBD-MTE method, which agree with each other within the error limits mentioned before. Our TEBD implemtation utilizes the ITensor library [25].
@@ III.4.1 Algorithm
In TEBD-MTE, the quantum many-body state $|\psi(t)\rangle$ of the electronic sub-system is described by a matrix product state (MPS). The combined state of the whole system $(|\psi_{\{x_{\ell},p_{\ell}\}}(t)\rangle,\{x_{\ell},~p_{\ell}\})$ is graphically represented in the bottom layer of Fig. 1. Figure 1 describes one graph element associated with a site $\ell$, which contains information about both the sub-systems, classical and quantum. Here, $A_{a_{\ell-1},a_{\ell}}^{\sigma_{\ell}}$ are the matrices of the MPS where $a_{\ell-1},a_{\ell}$ are the bond indices, and $\sigma_{\ell}$ is the physical index [85] and $x_{\ell}$ and $p_{\ell}$ represent the position and momentum coordinates of the classical phonon on site $\ell$.
@@
Figure 4: Dependence of the TEBD-MTE method on the discarded weight $\delta$ and the time step $\Delta t$ in an interacting system ($L=14$, $V=$2\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, $\gamma=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$, $W=0$). We plot the deviation $\delta\mathcal{O}_{\mathrm{CDW}}$ when 4 varying $\delta$ at a fixed time step $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$, comparing with a reference set using $\delta=10^{-8}$ and 4 varying $\Delta t$ at fixed cutoff $\delta=10^{-8}$}, comparing with a reference set using $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$}. The results are averaged over $N_{\mathrm{traj}}=4000$ trajectories.
@@
As we know [85, 75], in the TEBD method, the error builds up with time, which depends on the time step in the Trotter-Suzuki decomposition and the truncation of the singular values during the singular value decomposition (SVD). The latter is related to the entanglement increase in a nonequilibrium problem and leads to an increase in the required resources. We can control the error of the TEBD step by decreasing the time step and the discarded weight.
@@ III.5 Benchmark
We envision several interesting applications of TEBD-MTE beyond the cases covered here. Generally, one of the most important applications of time-dependent MPS methods is the calculation of equilibrium spectral functions and other dynamical properties. These are usually cases with a mild entanglement increase, even at finite temperature. While the regime of intermediate to large values of $\omega_{0}$ have been covered with full-quantum MPS methods [45, 47], the adiabatic regime of optical phonons with interacting electrons is mostly unexplored with MPS methods, which also applies to acoustic phonons. In the long run, using hybrid MPS-trajectory-based methods in DMFT might be an interesting direction as well. In order to illustrate that TEBD-MTE is well behaved concerning its convergence properties on larger systems, we present further results in Appendix 15.
@@ IV Stability of an Anderson insulator
In this section, we discuss the case of non-interacting electrons ($V=0$) with finite disorder. When the electronic sub-system is isolated from the phonons, the system exhibits Anderson localization for any finite disorder strength in one dimension [53]. In this section we study the stability of the Anderson localization in the presence of classical optical phonons.
Our expectation is that for small $\gamma$ and $\omega_{0}$, MTE will accurately reproduce the electronic observables even of the original quantum model. However, we will go beyond the regime of $\unit\gamma\ll\unit{}$ addressing the question of the stability of Anderson laocalization against a classical bath for arbitrary values of $\gamma$, regardless of whether the quantum version is accurately described or not. The same philosophy will apply in Sec. V. This section serves to set the stage for the analysis of systems with electronic interactions; so long as $V=0$, no many-body treatment of the electronic sector is necessary and instead, we use single-particle orbitals in the MTE implementation as in [97]. We first study the spreading of a single electron and then move on to the decay of the CDW initial state.
@@ IV.1 Spreading of a single electron
Figure 7: Time-evolved density profile $\langle\hat{n}_{\ell}(t)\rangle$ of a particle initially ($t=0$) localized on a single lattice site $\ell_{0}=50$ on a lattice with $L=100$ sites and with disorder ($W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$}). We show the 7 uncoupled case $\gamma=0$, exhibiting Anderson localization, and 7 a system with a large electron–phonon coupling $\unit\gamma=\unit{}$. Further parameters are $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$ and the data was generated using the Lanczos-MTE routine with $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$, $\epsilon=10^{-9}/t_{f}$, and $N_{\mathrm{dis}}=5000$ realizations for $N_{\mathrm{traj}}=200$ trajectories each.
@@
Figure 8: Spreading of a single particle in the disordered system coupled to classical phonons. For a few values of $\gamma$ we show in 8 the MSD for the particle as a function of time and in the inset the time-dependent dynamical exponent $\bar{z}(t)$, averaged over a time window of $t^{*}=$800\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$ (see Eq. 30). In 8 the extracted estimates for the dynamical exponent $z$ is shown. The model parameters are $L=100$, $\omega_{0}=$0.1\text{\,}\mathrm{\mathit{t}_{0}}$\$, and $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$ and the particle is initially localized at lattice site $\ell_{0}=50$. The data are generated using an MTE routine using exact diagonalization in the single-particle picture to propagate the quantum-state with $N_{\mathrm{dis}}=2000$ realizations for $N_{\mathrm{traj}}=200$ trajectories each and $\Delta t=$0.01\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$.
@@
Figure 9: Similar setup to Fig. 8 with $\unit\gamma=\unit{}$ held fixed while $W$ is varied. We show the dependency of the dynamical exponent $z$ on disorder strength obtained from MTE. Our estimate of the dynamical exponent $z$ is averaged over a large time window $t^{*}=$800\text{\,}{\mathrm{\mathit{t}_{0}}}^{-1}$\$. The solid symbols are data from a full quantum-treatment of electrons coupled to dispersive hard-core bosons for $\unit\omega_{0}=\gamma=\unit{}$ from [82], with a hard-core boson hopping amplitude $t_{b}=$0.4\text{\,}\mathrm{\mathit{t}_{0}}$\$.
Anderson localization implies no transport at all in one dimension [24], so any persistent spreading of a particle can be interpreted as a breakdown of Anderson localization. In Fig. 7 we show how a particle on a lattice of $L=100$ sites and disorder strength $W=$8\text{\,}\mathrm{\mathit{t}_{0}}$\$ spreads out in the system over time. The particle is initially localized at site $\ell_{0}=50$. For $\gamma=0$ we clearly see that the particle stays localized as expected. In contrast, when $\unit\gamma=\unit{}$, the particle continuously spreads out with a characteristic cone whose shape is related to the specific transport properties of the system.
@@
Interestingly, the qualitative and quantitative dependence of $z$ on $W$ is very similar to that obtained for $\unit\omega_{0}=\gamma=\unit{}$ obtained in [82] from a numerically exact approach studying an electronic system coupled to dispersive hard-core bosons (see the circles in Fig. 9).
In summary, we observe delocalization and subdiffusive electronic transport due to the coupling to classical phonons. Going to more electrons and interacting system is expected to accelerate that delocalization which we will capture by studying the decay of $\mathcal{O}_{\text{CDW}}$.